OpenStudy (anonymous):

Solve the equation by graphing x^2 - 5x - 24 = 0 A. {4, 6} B. {-2, 12} C. {-4, 6} D. {-3, 8}

OpenStudy (anonymous):

\[ x^2 - 5x - 24 = 0 \rightarrow x^2 - 8x+3x - 24 = 0\] \[ \rightarrow x(x - 8)+3(x - 8) = 0\] \[ \rightarrow (x - 8)(x + 3) = 0\] therefore either (x - 8) =0 or (x + 3) = 0 i.e. x=8 or x=-3 x={-3, 8} Hence option D. {-3, 8} is the correct answer. @AntoinetteWillman

OpenStudy (anonymous):